A ball is thrown straight up from the edge of the roof of a building of height 60.0m. A second ball is dropped from the roof 2s later. What must be the initial speed of the first ball if both are to hit the ground at the same time?
Answer #1

Time taken by second ball =

S= ut + 1/2 at^2 put u = 0

S = 1/2 × 9.8 t^2

60 ×2 /9.8 = t^2

t= 3.5s

Time taken by first ball = 3.5 + 2= 5.5s

-60 = u x 5.5 - 9.8/2 × 5.5 x5.5

u= 16m/s